OpenStudy (anonymous):

Ellason wants to solve the following system using the elimination method: 5x + 2y = 30 x + y = 8 What number should the equation x + y = 8 be multiplied by to eliminate y? 2 −2 5 −5

OpenStudy (anonymous):

Since the value of y in the first equation is 2, we can multiply the opposite of 2 to the second equation to eliminate y. So what is the opposite of 2?

OpenStudy (anonymous):

-2

OpenStudy (imtiaz7):

by 2 and then subtract them the y'll be eliminated

OpenStudy (anonymous):

Yes, so that is your answer. :)

OpenStudy (anonymous):

can you help me with another one @AnswerMyQuestions

OpenStudy (anonymous):

Sure. :)

OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? 2x + y = 20 4x − 2y = 40 (10, 0) (0, 10) (4, 2) (10, −10)

OpenStudy (imtiaz7):